how much pure acid should be mixed with 7 gallons of 50% acid solution in order to get an 80% acid solution?
Answer by KnightOwlTutor(293)

Pure acid is 100%
Convert % into decimal form 100/100
50%=50/100
80%=80/100
X=# of gallons of pure acid
1*x+7(.5)=(x+7)(.8)
Use the distributive property

x+3.5=.8x+5.6
Subtract .8x from both sides
.2x+3.5=5.6
Subtract 3.5 from both sides

.2x=2.1
Divide both sides by .2
x=10.5 gallons of pure acid
Let's double check our answer

10.5+3.5=(10.5+7)(.8)
14=14